A uniform electric field exists in the region between two oppositely charged plane parallel plates. A proton is released from rest at the surface of the positively charged plate and strikes the surface of the opposite plate, 1.60 cm distant from the first, in a time interval of 3.20 x 10-6 s. 

(a) Find the magnitude of the electric field.

 (b) Find the speed of the proton when it strikes the negatively charged plate.




1
Expert's answer
2022-01-13T09:14:52-0500

Solution

Time interval

t=3.20 x 10"^{-6}" s

Distance between plates

d=1.6cm

Charge of proton

q=1.6"\times10^{-19}C"

Mass of proton

M=1.67"\times10^{-27}Kg"


a) so magnitude of the electric field

"E=\frac{2dM}{qt^2}\\E=\frac{2\times0.016\times1.67\times10^{-27}}{1.6\times10^{-19}(3.20\times10^{-6})^2}"

E=32.62V/m

b) now speed

"v=\frac{qEt}{m}\\=\frac{1.6\times10^{-19}\times32.62\times3.2\times10^{-6}}{1.67\times10^{-27}}\\=10000m/sec."